OpenStudy (asylum15):

Find the following indefinite integral : integral (1+x^2)sin(x^3+3x)dx

OpenStudy (asylum15):

\[\int\limits(1+x^2)\sin(x^3+3x)dx\]

OpenStudy (asylum15):

Hey man, I'm a biomed engineer, just revising some integration. Listen, I know the basics here, but when I got stuck in the question, can I ask you about where I got stuck?

OpenStudy (asylum15):

Heres how far i got: Let u = x^3 + 3x du/dx = 3x^2+3 => dx = du/3x^2+3 From here?

terenzreignz (terenzreignz):

Then just replace dx in your integral with du/3x² + 3 Better yet, take notice of the fact that a constant can be factored out of 3x² + 3

OpenStudy (asylum15):

I got \[\int\limits \sin(u) . (1+x^2) . du/3x^2+3\]

terenzreignz (terenzreignz):

Hang on, easier to see it like this: \[\int\limits\limits(1+x^2)\sin(x^3+3x)\frac{du}{3 + 3x^{2}}\]

terenzreignz (terenzreignz):

So, the denominator of 3 + 3x², 3 can be factored out, right?

OpenStudy (asylum15):

1/3 ?

OpenStudy (asylum15):

Outside the integral

terenzreignz (terenzreignz):

That's right ; \[\int\limits\limits\limits\sin(x^3+3x)\frac{(1 + x^{2})du}{3(1 + x^{2})}\] Consequently, as you said: \[\frac{1}{3}\int\limits_{}^{} \sin(x^{3}+3x)du\] Carry on? :D

OpenStudy (asylum15):

So \[\frac{ 1 }{ 3 }\int\limits \sin(u) . du ?\]

terenzreignz (terenzreignz):

Precisely :D

OpenStudy (asylum15):

See, I'm excellent at method, its SEEING manipulations I'm weak at. Comes with practice?

terenzreignz (terenzreignz):

Indeed it does :D

OpenStudy (asylum15):

Ok, do you mind if I ask you one other quick thing? Then I'm sorted.

terenzreignz (terenzreignz):

Go ahead

OpenStudy (asylum15):

\[\int\limits \frac{ 7dx }{ (x-4)(x+2) }\]

OpenStudy (asylum15):

Partial fraction?

terenzreignz (terenzreignz):

It would certainly seem like that's the way to go. You know where to begin?

OpenStudy (asylum15):

I'm thinking bring the 7 outside the integral first?

terenzreignz (terenzreignz):

You could do that: \[7\int\limits_{}^{}\frac{dx}{(x-4)(x+2)}\] And then?

OpenStudy (asylum15):

Do partial fractions of 1/(x-4) + 1/(x+2)

OpenStudy (asylum15):

Is there another method? You can't do a substitution, and LIATE doesnt work?

terenzreignz (terenzreignz):

Substitution doesn't seem possible, and I don't know what LIATE is o.O And I'm afraid it's no simple matter of separating the fraction with 1 as the numerator on both sides. Don't worry though, it may not be THAT simple, but it IS rather simple :) Your factors (x + 2) and (x - 4) are both linear, or or degree 1. Then all we know about their respective numerators is that they'll be of one LESS degree, and therefore their numerators are just constants. Let's call them A and B. \[\frac{1}{(x-4)(x+2)}=\frac{A}{x-4}+\frac{B}{x+2}\] Can you do it from here?

OpenStudy (asylum15):

Yes, I can do that. I'm just curious to how you will then integrate? By the way, LIATE = Logs, inverse trig, algebra, trig, exponential, When looking at a question, I see if it has any of these, to decide what to do.

terenzreignz (terenzreignz):

Well, integration distributes over a sum, and integrating a constant over a linear function is simple, to illustrate

terenzreignz (terenzreignz):

\[\int\limits_{}^{}\frac{k}{ax+b}dx=\frac{k}{a}\ln|ax + b| +C\]

OpenStudy (asylum15):

Yeah, was going to mention log would come into it. Will we end up with a LogA - Log B?

terenzreignz (terenzreignz):

No. Let's carry on from \[\frac{1}{(x-4)(x+2)}=\frac{A}{x-4}+\frac{B}{x+2}\] \[=\frac{A(x+2)+B(x-4)}{(x-4)(x+2)}\]

OpenStudy (asylum15):

Ok

terenzreignz (terenzreignz):

\[=\frac{(A+B)x+2A-4B}{(x-4)(x+2)}=\frac{1}{(x-4)(x+2)}\] Is it clearer from here?

OpenStudy (asylum15):

Yes

terenzreignz (terenzreignz):

Well, can you do it from there?

OpenStudy (asylum15):

Pretty sure i can! :D

OpenStudy (asylum15):

Thanks, mate :)

terenzreignz (terenzreignz):

No problem